# Complete Neural Networks for Complete Euclidean Graphs

### v3

@@ Introduction
Dym and Maron (2021) proved that the well-known Tensor Field Network (Thomas et al. 2018) invariant architecture is universal, but the construction in their proof requires arbitrarily high-order representations of the rotation group. Similarly, universality can be obtained using high-order representations of the permutation group (Lim et al. 2022). However, before this work, it was not known whether the same theoretical guarantees can be achieved by realistic point-cloud architectures that use low-dimensional representations, and whose complexity has a mild polynomial dependency on the data dimension. In the words of (Pozdnyakov and Ceriotti 2022): ”…provably universal equivariant frameworks are such in the limit in which they generate high-order correlations… It is an interesting, and open, question whether a given order suffices to guarantee complete resolving power.” (p. 6). We note that it is known that separation of point clouds in polynomial time in $n$ is possible, assuming that $d$ is fixed (e.g., $d=3$) (Arvind and Rattan 2014; Dym and Kovalsky 2019; Kurlin 2022). What still remains to be established is whether separation is achievable for common invariant machine learning models, and more generally, whether separation can be achieved by computing a continuous invariant feature that is piecewise differentiable.
@@
For general fixed $d$, there do exist algorithms that can separate point clouds up to equivalence in polynomial time, but they do not seem to lend themselves directly to neural architectures. In (Kurlin 2022; Widdowson and Kurlin 2023) complete tests are described, but they represent each point cloud as a ‘multiset of multisets’ rather than as a vector as we do, and so are not suitable for gradient-descent-based learning. Efficient tests for equivalence of Euclidean graphs were described by Brass and Knauer (2000); Arvind and Rattan (2014), but they compute features that do not depend continuously on the point cloud.
@@
Widdowson and Kurlin (2022) suggest a method for distinguishing almost every point clouds up to equivalence, similar to our result here on $1$-EWL. Similarly, efficient separation/universality can also be obtained for point clouds with distinct principal axes (Puny et al. 2021; Kurlin 2022). Another setting in which universality is easier to obtain is when only rigid symmetries are considered and permutation symmetries are ignored (Wang et al. 2022; Villar et al. 2021; Satorras, Hoogeboom, and Welling 2021). All these results do not provide universality for all point clouds, with respect to the joint action of permutations and rigid motions.
@@
$\Psi_{sort}(x_{1},\ldots,x_{n})=\mathrm{sort}(x_{1},\ldots,x_{n})\quad$ (6)
@@
$\text{{Embed}}_{\theta}(x_{1},\ldots,x_{n})=\langle b_{j},\;\Psi\left(a_{j}^{T}x_{1}\ldots,a_{j}^{T}x_{n}\right)\rangle,$ (8)

### v4

@@ Introduction
Dym and Maron (2021) proved that the well-known Tensor Field Network (Thomas et al. 2018) invariant architecture is universal, but the construction in their proof requires arbitrarily high-order representations of the rotation group. Similarly, universality can be obtained using high-order representations of the permutation group (Lim et al. 2022). However, before this work, it was not known whether the same theoretical guarantees can be achieved by realistic point-cloud architectures that use low-dimensional representations, and whose complexity has a mild polynomial dependency on the data dimension. In the words of (Pozdnyakov and Ceriotti 2022): ”…provably universal equivariant frameworks are such in the limit in which they generate high-order correlations… It is an interesting, and open, question whether a given order suffices to guarantee complete resolving power.” (p. 6). We note that it is known that separation of point clouds in polynomial time in $n$ is possible, assuming that $d$ is fixed (e.g., $d=3$) (Arvind and Rattan 2016; Dym and Kovalsky 2019; Kurlin 2024). What still remains to be established is whether separation is achievable for common invariant machine learning models, and more generally, whether separation can be achieved by computing a continuous invariant feature that is piecewise differentiable.
@@
For general fixed $d$, there do exist continuous algorithms that can separate point clouds up to equivalence in polynomial time, but they do not seem to lend themselves directly to neural architectures. In (Kurlin 2024; Widdowson and Kurlin 2023) a polynomial-time algorithm is introduced for computing invariants that completely determine a point cloud that is Lipschitz-continuous, yet these invariants are represented as a ‘multi-set of multi-sets’ and thus do not allow for gradient-descent-based optimization. Our approach is continuous and represents the point cloud as a vector in real space, which allows for back-propagation. Efficient tests for equivalence of Euclidean graphs were described by Brass and Knauer (2000); Arvind and Rattan (2016), but they compute features that do not depend continuously on the point cloud.
@@
Widdowson and Kurlin (2022) suggest a method for distinguishing almost every point clouds up to equivalence, similar to our result here on $1$-EWL. Similarly, efficient separation/universality can also be obtained for point clouds with distinct principal axes (Puny et al. 2021; Kurlin 2024). Another setting in which universality is easier to obtain is when only rigid symmetries are considered and permutation symmetries are ignored (Wang et al. 2022; Villar et al. 2021; Satorras, Hoogeboom, and Welling 2021). All these results do not provide universality for all point clouds, with respect to the joint action of permutations and rigid motions.